# Reva Schwartz1    Carina Westling2    Morgan Briggs3    Marzieh Fadaee4    Isar Nejadgholi5    Matthew Holmes6    Fariza Rashid7    Maya Carlyle8    Afaf Taïk9    Kyra Wilson10    Peter Douglas11    Theodora Skeadas12    Gabriella Waters1,13    Rumman Chowdhury14    Thiago Lacerda3

### 03 Mar 2026 v2

@@ Abstract
This paper proposes CIRCLE, a six-stage, lifecycle-based framework to bridge the "reality gap" between model-centric performance metrics and AI’s materialized outcomes in deployment. While existing frameworks like MLOps focus on system stability and benchmarks measure abstract capabilities, decision-makers outside the AI stack lack systematic evidence about the behavior of AI technologies under real-world user variability and constraints. CIRCLE operationalizes the "Validation" phase of TEVV (Test, Evaluation, Verification, and Validation) by formalizing the translation of stakeholder concerns outside the stack into measurable signals. Unlike participatory design which often remains localized, or algorithmic audits which are often retrospective, CIRCLE provides a structured, prospective protocol for linking context-sensitive qualitative insights to scalable quantitative metrics. By integrating methods such as field testing, red teaming, and longitudinal studies into a coordinated pipeline, CIRCLE produces "systematic knowledge"—evidence that is comparable across sites yet sensitive to local context. This can enable governance based on materialized downstream effects rather than theoretical capabilities.
1 A shift in evaluation goals
@@
As adoption of generative AI continues to rise, policymakers, organizations, and the public are increasingly asking not just what these tools can do, but whether they can create meaningful value and work safely in the real world [22, 40]. Organizations want to identify viable AI use cases in advance and convert pilot deployments into durable successes, while the public expects that AI will work as advertised and not expose them, their families, or their communities to unsafe or harmful outcomes [61, 62, 49, 40]. AI evaluation is the applicable discipline for answering these questions, yet the current ecosystem’s focus inside the AI stack leaves it unable to accommodate the diversity of purposes and settings in which AI is being adopted and used [29]. The lifecycle described in this paper aims to address this gap by formalizing methods, tools, and processes that can produce systematic knowledge to support stakeholders beyond the AI development stack.
@@
Like other forms of technology, AI can contribute to long-term and broad-scale impacts that reshape human behavior, organizational structures, and societal processes over time [57]. These higher-order effects underscore the importance of evaluating AI from a real-world and purpose-driven lens. The dominant evaluation ecosystem measures AI’s primary effects-the immediate outputs of a model-under isolated settings. Secondary effects are the near term impacts those outputs produce in everyday life. For example, “AI psychosis” and other psychological effects associated with how users may over-rely on chatbots do not result from any single output, but from the cumulative interaction patterns between people and AI systems over time [6, 42, 27] .
@@
As AI is embedded into workplace operations, secondary effects can manifest as shifts in day‑to‑day workflows, redefinitions of job roles, and changes in how teams coordinate and make decisions [67, 66, 3, 16]. Over longer horizons, such changes can accumulate into tertiary effects on productivity, profitability, and overall return on investment (ROI) [1, 43]. AI’s ROI is currently estimated using self-reported data from surveys with executives [50]. These measures are more likely to capture executive expectations, narratives, and strategic signaling than realized operational changes, and are hard to corroborate without direct measures of deployed AI and its associated impacts [17, 11]. With AI’s broad‑scale societal shifts largely invisible to the current evaluation ecosystem, policymakers and organizations have limited insight to inform decisions about whether and how to deploy these systems in everyday settings.
@@
To address the gap between static model benchmarks and dynamic real-world outcomes, this paper presents CIRCLE not merely as a descriptive heuristic, but as a formal lifecycle framework that shifts the scope of AI assessment. While current evaluations often treat contextual factors as noise to be eliminated [52, 8, 63], CIRCLE treats it as the primary signal for understanding system behavior in the real-world. By focusing only on inside-the-stack activities and abstract model capabilities, many existing approaches both fail to support deployment stakeholders and introduce threats to core forms of measurement validity–such as ecological, construct, and consequential validity [58, 41]. This can undermine the credibility of inferences about how models will behave once they are integrated into real-world contexts and workflows [52, 34]. CIRCLE instead contributes a structured methodological protocol for generating systematic knowledge about what materializes in AI deployment. Specifically, it delivers:
@@
  • An Integrated Lifecycle Model: A six-stage architecture that unifies currently isolated evaluation activities—such as stakeholder elicitation (Stage 1), red teaming (Stage 3), and longitudinal monitoring (Stage 6)—into a coherent, iterative feedback loop. This brings a traceable pipeline to current ad-hoc testing paradigms where every stage generates its own documented output, feeding directly into the next stage to build a continuous record of evaluation.
@@
  • A Construct Operationalization Schema: A specific method for translating the concepts that matter to stakeholders outside the AI stack (e.g., "over-reliance" or "cognitive offloading") into observable behavioral indicators and quantitative metrics. This process bridges the gap between qualitative socio-technical concerns and quantitative measurement and results in a context brief that serves as the foundation for subsequent stages of the evaluation lifecycle.
@@ 2 The CIRCLE framework
CIRCLE—short for Contextualize, Identify, Represent, Compare, Learn, and Extend— is a six-stage lifecycle framework that applies three specific methodological shifts to address specific unresolved gaps in the current evaluation ecosystem:
@@
    Treating real-world heterogeneity as signal instead of noise: Model-centric benchmarking typically measures abstract capability under optimized conditions, treating user heterogeneity as noise to be eliminated [18, 63, 65, 9, 26]. In contrast, CIRCLE treats the variability in how people interpret and adapt to AI—as the primary signal of operational success. This can reveal the gap between technical accuracy metrics and actual deployment outcomes.
@@
    Bringing a socio-technical frame to measurement: Benchmarking assesses AI via script- based in silico testing and MLOps frameworks focus on technical stability, latency, and data drift. CIRCLE evaluates AI through the lens of people interacting with systems in real workflows to capture materialized outcomes. This extends measurement towards higher-order effects—such as cognitive offloading and long-run productivity—that technical telemetry alone cannot detect.
@@
    Moving beyond evaluating AI "one context at a time" to aggregate knowledge across settings: While participatory design generates rich local insights, it often produces fragmented signals that are difficult to compare across sites. CIRCLE combines the contextual depth of participatory methods with the rigorous, repeatable structure of industrial testing. This allows organizations to aggregate evidence and compare disparate systems using shared constructs.
@@
Figure 1 presents the overall structure of the CIRCLE lifecycle, emphasizing its iterative nature, the flow of contextualized evaluation evidence across stages, and the role of secondary and tertiary effects as evaluation targets. Each lifecycle stage in the figure also includes a brief description of its primary output.
Figure 2 shows how the lifecycle can be used to take a stakeholder-specified concern, systematize it into a construct, translate it into observable behavior, and link it to measures that can reveal higher‑order effects.
CIRCLE can operationalize the "Validation" phase of the Test, Evaluation, Verification, and Validation TEVV- framework [44] by shifting the focus from whether models match prescribed criteria to aggregate descriptive reporting of what happens when people use AI in realistic conditions. Its main contribution is a construct-centered TEVV process that a) starts from concerns and contextual factors defined by deployment-level stakeholders rather than available datasets (b) systematizes those concerns into measurable constructs rather than using convenient proxies, and (c) coordinates evaluation methods within a single, traceable pipeline that can be run in situ and at scale rather than as one‑off, small‑sample or purely lab-based tests. By tying every metric back to a named construct and a stakeholder-relevant outcome, CIRCLE produces evidence that is not just technically accurate, but operationally valid for decision-making.
To make each stage more concrete, this section uses an EdTech (educational technology) example in which teachers and students may over-rely on AI chatbots, which could result in shifts in reliance and instructional quality, over time. Throughout the remainder of this section, stage‑specific figures highlight the primary activities at each lifecycle stage by emphasizing the focal stage and greying out the others in the sequence.
@@
Identify teEvaluation design and planning stage Translates stakeholder-defined concepts of interest into test-ready methods and data‑collection protocols. What forms of evidence need to be produced in order to judge whether system outcomes can match stakeholder expectations in this setting?
@@ 2.1 Stages of the lifecycle
The CIRCLE lifecycle structures activities to define what matters in a setting, and then designs and runs context‑aware tests, interprets results for decision‑makers, and feeds insights back into ongoing monitoring. Each stage produces a formal output deliverable that becomes the actionable input and “to‑do” list for the next stage, creating a traceable chain of evaluation work products. Table 1 summarizes the key concepts and guiding questions associated with each stage.
A key differentiator in the CIRCLE framework is how it leverages people in a more contextually centered way than most current AI evaluations, where users are typically involved either to generate training data or rate model outputs in isolation. Instead of treating people as one‑off labelers, CIRCLE engages them as situated experts on how system behavior aligns with their goals, constraints, and local norms, and pairs their accounts with interaction transcripts to reveal patterns that matter for real‑world use. These human‑centered collections can run alongside large‑scale automated test runs so that findings generalize beyond a narrow test group and evaluation mode. Each element of CIRCLE corresponds to specific activities in a lifecycle stage, linking stakeholder‑defined concepts of interest to practical evaluation methods and evidence.
Patterned on Chouldechova et al.’s shared standard for valid measurement of generative AI systems [13], the CIRCLE lifecycle shifts their core categories from Concepts to Concepts and Questions, Instance to Processes, Population to Who and Where, and Amounts to Outcomes. The lifecycle focuses on practical real-world context, materialized impacts, and stakeholder input, broadening who defines evaluation questions, their scope, and what counts as evidence. The full descriptive information for each CIRCLE lifecycle stage is provided in Table LABEL:tab:frame-lifecycle in the Appendix.
Real-world evaluation activities extend beyond the traditional AI stack environment to settings where people decide whether and how to leverage AI. This broader remit demands different organizational capabilities and resourcing compared to conventional model-testing. For example, organizations must be able to support stakeholder engagement, construct articulation, instrument design, and quasi‑experimental testing with large and sufficiently diverse populations across various settings. The ability to conduct and sustain these activities over time may be out of reach for many individual teams acting alone, underscoring the need for shared methods, tools, and support structures that lower the barrier to participation in robust, context‑sensitive evaluation.
@@ 2.1.1 Stage 1: Context specification
Since situational factors are experienced differently across roles and communities, the lifecycle’s first stage builds contextual awareness by eliciting what matters most to stakeholders in a given setting [4, 37, 64, 63, 35, 14]. Without this process, the follow-on stages lack a grounded basis for interpreting results, making it difficult to judge whether observed outcomes meaningfully reflect the concerns, priorities, and constraints of the people affected.
Stage 1 activities center on eliciting information from key stakeholders about the AI system’s purpose, the operational constraints that shape its behavior, and the stakes for those directly or indirectly affected by its deployment [53]. In the education example, school leaders, educators, and communities can jointly specify their priorities, concerns, and expectations for AI chatbot deployment, along with relevant regulatory and ethical boundaries.
Elicitation activities can include in-person workshops, interviews, and process mapping, or they can be conducted virtually through asynchronous exercises and other engagements, offering flexibility around organizational time and resource constraints. These techniques gather both explicit knowledge (documented policies, procedures, and formal requirements) and the tacit knowledge (experiential insights, practical judgment, and uncodified understanding) necessary to understand the deployment context and how organizations actually operate [25].
These qualitatively defined priorities can be augmented by automated methods, such as LLM‑based review of documented curricula, policies, and community materials, to surface salient risks and considerations that stakeholders may overlook [23, 15, 36]. Combining human‑centered elicitation with automated techniques offers a fuller picture of both tacit and explicit knowledge, aligning with work that frames AI and humans as complementary partners in knowledge work. This approach also makes it feasible to extend contextual analysis across many settings that match the prospective use case [30].
The collected information is used to build up a “construct” – or a clear, named idea that represents the specific risk, behavior, or value that the stakeholders care about, along with a description of what the construct means in practice, when it shows up, and what counts as acceptable or concerning levels. Constructs of interest in the EdTech example might include student and teacher over‑reliance on chatbot outputs, cognitive offloading, and preserving human judgment in the classroom. Evaluation designers translate contextual information into constructs by judging which stakeholder concerns can be addressed in a given setting and meaningfully operationalized as measurable indicators. Figure 3 shows how Stage 1 activities focus on eliciting and systematizing stakeholder‑specified concerns.
The output of this stage is a “context brief” that feeds into Stage 2, where constructs are translated into concrete evaluation parameters.
@@ 2.1.2 Stage 2: Evaluation design and planning
In Stage 1, the stakeholders in the EdTech example raise a concern that “teachers and students might defer their own decisions and judgments to an educational chatbot.” In Stage 2, this concern is formalized as a construct such as “changes in content review and deferring of decisions and judgments,” making it specific enough to guide observation and measurement.
Figure 4: Figure 4. Stage 2 of the lifecycle in the EdTech example: designing test activities to assess the presence of the defined construct.
The next step is to design evaluation methods and scenarios able to instantiate the contextual factors that stakeholders consider most relevant. This process contrasts with model-centric evaluations which wall off contextual details by tightly controlling test environments, static datasets, and canned prompts. Stakeholders beyond the stack require contextual detail to make sense of evaluation results in setting and support informed deployment decisions.
Figure 5 illustrates how evaluation methods can be selected and combined as design choices in Stage 2, highlighting tradeoffs between control and contextual richness and how those choices shape what forms of evidence can be observed.
Figure 5: Evaluation methods as design choices in Stage 2 of the CIRCLE lifecycle. Methods trade off control and contextual richness, shaping what forms of evidence and downstream effects can be observed. Regions illustrate classes of methods that may be selected and combined during this stage.
At the primary end of the continuum is AI benchmarking, which focuses on whether models produce “correct” or expected outputs and yields metrics such as accuracy, efficiency, and energy use. Methods such as usability studies begin to surface AI’s secondary effects. By examining whether people can successfully navigate AI interfaces to complete tasks, these approaches help to reveal user friction, confusion, or drop‑off points that simple benchmarks cannot detect.
Context‑sensitive methods such as red‑teaming, bias bounties, and field tests evaluate system behavior during live interactions with people under realistic or adversarial conditions [38, 31]. At the tertiary end, longitudinal studies can be combined with other real-world signals like incident tracking to monitor cumulative downstream effects that no single snapshot can capture, such as how innovation benefits for organizations, sectors, and the broader economy evolve over time.
Stage 2 activities drive decisions about which methods along this continuum are best suited to evaluate the construct and generate insights that stakeholders can apply in their own setting. This stage also defines the tasks and scenarios for capturing the required evidence from participants at scale. For example, different designs can test whether privacy guardrails hold in practice, how often high‑stakes errors occur in a given setting, and whether multilingual capabilities genuinely support an intended population [28].
In parallel, large‑scale automated test runs are designed to evaluate the same constructs under controlled conditions by stress‑testing models on synthetic or replayed interaction logs that mirror the scenarios used in human trials. These automated runs extend coverage across edge cases and input variations that are impractical to reach solely through human studies.

### 18 Mar 2026 v3

@@ Abstract
This paper proposes CIRCLE, a six-stage, lifecycle-based framework to bridge the "reality gap" between model-centric performance metrics and AI system outcomes in deployment. Current approaches such as MLOps frameworks and AI model benchmarks offer detailed insights into system stability and model capabilities, but they do not provide decision-makers outside the AI stack with systematic evidence of how these systems actually behave in real‑world contexts or affect their organizations over time. CIRCLE operationalizes the "Validation" phase of TEVV (Test, Evaluation, Verification, and Validation) by translating priorities of stakeholders outside the stack into measurable signals. Unlike participatory design which often remains localized, or algorithmic audits which are often retrospective, CIRCLE provides a structured, prospective protocol for linking context-sensitive qualitative insights to scalable quantitative metrics. By integrating methods such as field testing, red teaming, and longitudinal studies into a coordinated pipeline, CIRCLE produces "systematic knowledge"—evidence that is comparable across sites yet sensitive to local context. This can enable governance based on materialized downstream effects rather than theoretical capabilities.
keywords:
real-world AI evaluation; socio-technical systems; testing, evaluation, verification, validation (TEVV)
1 Introduction
@@
As adoption of generative AI continues to rise, policymakers, organizations, and the public are increasingly asking not just what these tools can do, but whether they can work safely and create meaningful value in the real-world [21, 39]. Organizations want to identify viable AI use cases in advance and convert pilot deployments into durable successes, while the public expects that AI will work as advertised and not expose them, their families, or their communities to unsafe or harmful outcomes [60, 61, 48, 39]. AI evaluation is the applicable discipline for answering these questions, yet the development‑centric lens of the current ecosystem is unable to accommodate the diversity of purposes and settings in which these technologies are being adopted and used [27]. The lifecycle described in this paper aims to address this gap by formalizing methods, tools, and processes that can produce systematic knowledge to support decision making beyond the AI development stack.
@@
Like other forms of technology, AI can contribute to long-term and broad-scale impacts that reshape human behavior, organizational structures, and societal processes over time [56]. These higher-order effects underscore the importance of evaluating AI from a real-world and purpose-driven lens. The dominant evaluation ecosystem focuses on AI’s primary effects—the immediate outputs of a model in isolated settings—rather than the secondary effects those outputs produce in everyday life. For example, “AI psychosis” and other psychological effects linked to users’ over‑reliance on chatbots do not result from any single output, but from cumulative interaction patterns between people and AI systems over time[6, 41, 26].
@@
As AI is embedded into workplace operations, secondary effects can manifest as shifts in day‑to‑day workflows, redefinitions of job roles, and changes in how teams coordinate and make decisions [66, 65, 3, 16]. Over longer horizons, such changes can accumulate into tertiary effects on productivity, profitability, and overall return on investment (ROI) [1, 42]. AI’s ROI is currently estimated using self-reported data from surveys with executives [49]. These measures are more likely to capture executive expectations, narratives, and strategic signaling than realized operational changes, and are difficult to corroborate without direct measures of deployed AI and its associated impacts [17, 11]. With AI’s broad‑scale societal shifts largely invisible to the current evaluation ecosystem, policymakers and organizations have limited insight to inform decisions about whether and how to deploy these systems in everyday settings.
@@
The CIRCLE lifecycle shifts the scope of evaluation to address the gap between static, model‑centric benchmarks and the dynamic higher-order effects posed by the use of AI across a variety of contexts. Current evaluations often treat such contextual factors as noise to be eliminated [51, 8, 62], which can threaten core forms of measurement validity—such as ecological, construct, and consequential validity [57, 40]—and undermine the credibility of inferences about how models will behave once deployed[51, 33]. The CIRCLE lifecycle leverages context as the primary signal of system behavior to generate systematic knowledge about how AI systems perform in real‑world deployments. Specifically, it delivers:
@@
  • An Integrated Lifecycle: A six-stage architecture that unifies currently isolated evaluation activities—such as stakeholder elicitation (Stage 1), red teaming (Stage 3), and longitudinal monitoring (Stage 6)—into a coherent, iterative feedback loop. This brings a traceable pipeline to current ad-hoc testing paradigms where every stage generates its own documented output, feeding directly into the next stage to build a continuous record of evaluation.
@@
  • A Construct Operationalization Schema: A method for translating the concepts that matter to stakeholders outside the AI stack (e.g., "over-reliance" or "cognitive offloading") into observable behavioral indicators and quantitative metrics. This process bridges the gap between qualitative socio-technical concerns and quantitative measurement–yielding a context brief that serves as the foundation for subsequent stages.
@@
Definitions for key AI systems terminology are listed in the “AI Deployment and System Terms” box below.
@@ 2 The CIRCLE framework
CIRCLE—short for Contextualize, Identify, Represent, Compare, Learn, and Extend— shifts evaluation methodology to address the following key gaps in the current evaluation ecosystem:
@@
    Treating real-world heterogeneity as signal instead of noise: Model-centric benchmarking typically measures abstract capability under optimized conditions, filtering out real-world variability[18, 62, 64, 9, 25]. CIRCLE instead treats differences in how people interpret and adapt to AI as a primary signal of operational success, revealing gaps between technical accuracy metrics and actual deployment outcomes.
@@
    Bringing a socio-technical frame to measurement: Benchmarking uses script-based in silico testing, and MLOps frameworks focus on technical stability, latency, and data drift. CIRCLE extends evaluation to AI’s higher-order effect by examining how interactions with AI systems in context trigger mechanisms in real workflows that produce downstream outcomes for people and organizations.
@@
    Moving beyond evaluating AI "one context at a time" to capture contextual knowledge at scale and across settings: Participatory and contextual methods generate rich local insights but often produce fragmented signals that are difficult to compare across sites. CIRCLE combines the contextual depth of participatory methods with the structured, repeatable discipline of industrial testing, allowing organizations to use shared constructs and metrics for comparing systems across settings.
@@
Figure 1 shows how contextualized evaluation evidence flows iteratively across the CIRCLE lifecycle stages, with a brief description of the primary output for each stage.
The lifecycle helps operationalize the “Validation” phase of the TEVV framework [43] by shifting focus from model compliance with preset criteria to what actually happens when people use AI in realistic conditions. Its main contribution is a construct-centered TEVV process that a) starts from stakeholder-articulated priorities about deployment–rather than available datasets (b) systematizes those priorities into measurable constructs–rather than using convenient proxies, and (c) coordinates evaluation methods within a single, traceable pipeline that can be run in situ and at scale–rather than as one‑off, small‑sample or purely lab-based tests. By tying every metric back to a named construct and a stakeholder-relevant outcome, CIRCLE produces evidence that is not just technically accurate, but operationally valid for deployment-level decision-making.
This section uses an EdTech vignette to illustrate how each stage operates across the full lifecycle. Figure 2 shows how stakeholder priorities are elicited, organized into constructs, and linked to observable behaviors and higher‑order effects for the EdTech vignette. In this example, stakeholders identify teacher and student over-reliance on AI chatbots as a key evaluation priority since it can lead to changes in instructional quality and learning over time. Stage‑specific figures highlight key activities by emphasizing the focal stage and greying out the others in the sequence.
@@
Identify Evaluation design and planning stage Translates stakeholder-defined concepts of interest into test-ready methods and data‑collection protocols. What forms of evidence need to be produced in order to judge whether system outcomes can match stakeholder expectations in this setting?
@@ 2.1 Stages of the lifecycle
The CIRCLE lifecycle defines what matters in a setting, runs context‑aware tests, interprets results for decision‑makers, and feeds insights into ongoing monitoring. Each stage produces a deliverable that becomes the next stage’s input, forming a traceable chain of evaluation work products. Table 1 summarizes the key concepts and guiding questions associated with each stage.
CIRCLE differentiates itself by centering people who test AI systems as situated experts rather than one‑off labelers of model outputs. It asks them how system behavior aligns with their local goals, constraints, and norms, then pairs their accounts with interaction transcripts to surface real‑world usage patterns tied to stakeholder priorities. These human‑centered collections can run alongside large‑scale automated tests so that findings generalize beyond a narrow test group or evaluation mode, with each lifecycle stage linking stakeholder‑defined concepts to concrete methods and evidence.
Patterned on Chouldechova et al.’s shared standard for valid measurement of generative AI systems [13], the CIRCLE lifecycle shifts their core categories from Concepts to Concepts and Questions; Instance to Processes; Population to Who and Where; and Amounts to Outcomes. It centers practical context, materialized impacts, and stakeholder input, broadening who defines evaluation questions, their scope, and what counts as evidence. Full descriptive details for each lifecycle stage appear in Table LABEL:tab:frame-lifecycle in the Appendix.
Real‑world evaluation reaches beyond the technical stack into the settings where people decide whether and how to use AI. This broader remit requires different organizational capabilities and resourcing than conventional model testing. For example, organizations must be able to support stakeholder engagement, construct articulation, instrument design, and quasi‑experimental testing with large and sufficiently diverse populations across various settings. The ability to conduct and sustain these activities over time may be out of reach for many individual teams acting alone, underscoring the need for shared methods, tools, and support structures that lower the barrier to participation in robust, context‑sensitive evaluation.
The “AI Evaluation and Testing Terms” box below lists definitions for core evaluation terms found in this paper.
@@ 2.1.1 Stage 1: Context specification
Since situational factors are experienced differently across roles and communities, the lifecycle’s first stage builds contextual awareness by eliciting what matters most to stakeholders in a given setting [4, 36, 63, 62, 34, 14]. This grounding step provides the basis for interpreting results in later stages and for judging whether observed outcomes truly reflect the concerns, priorities, and constraints of stakeholders closest to the deployment context. Elicitation can happen through in‑person workshops, interviews, and process mapping, or virtually via asynchronous exercises, offering flexibility around time and resource constraints. These techniques surface both explicit knowledge (policies, procedures, formal requirements) and tacit knowledge (experiential insights, practical judgment, uncodified understanding) needed to grasp the deployment context and how organizations actually operate [24].
Stakeholder-elicited priorities can be augmented by automated methods, such as LLM‑based review of documented curricula, policies, and community materials, to surface considerations that stakeholders may overlook [22, 15, 35]. Combining human‑centered elicitation with automated techniques offers a more comprehensive picture of both tacit and explicit knowledge, aligning with work that frames AI and humans as complementary partners in knowledge work. This approach also makes it feasible to extend contextual analysis across many settings that match the prospective use case [29].
In the education example, school leaders, educators, and communities begin by specifying their priorities, concerns, and expectations for the classroom chatbot. Evaluators then situate these priorities within supporting information about the deployment context—the system’s purpose, operational constraints, regulatory and ethical boundaries, and the stakes for those affected. This collected information is then synthesized into an evaluation construct: a clear, named idea representing the specific risk, behavior, or value that stakeholders need evidence about to support their deployment decisions.. For example, identification of an "over‑reliance" construct might cover cognitive offloading and preserving human judgment in the classroom. This can be articulated as “teachers and students might defer their own decisions and judgments to an educational chatbot.” Figure 3 shows how Stage 1 elicits and systematizes these stakeholder priorities, producing a “context brief” that feeds into Stage 2.
@@ 2.1.2 Stage 2: Evaluation design and planning
Figure 4: Stage 2 of the lifecycle in the EdTech example: designing test activities to assess the presence of the defined construct.
Stage 2 formalizes the construct laid out in the context brief so it is specific enough to guide observation and measurement - for example “changes in content review and deferring of decisions and judgments”. Evaluators then design evaluation methods and scenarios that make the construct measurable and observable within the relevant context (e.g., the classroom setting). This foregrounding of context-specific factors contrasts with model‑centric evaluations, which typically use static datasets, canned prompts, and tightly controlled test environments. Relevant contextual detail helps stakeholders beyond the stack make sense of evaluation results in their own settings to support informed deployment decisions. Figure 4 show how Stage 2 activities operationalize these design choices for the EdTech example.
Figure 5 shows a continuum of evaluation methods that can be selected and combined to deploy the Stage 2 test scenarios, highlighting tradeoffs between pre-scripted, in‑silico tests and context‑rich, real‑world methods, and how those choices shape what forms of evidence can be observed.
Figure 5: Evaluation methods as design choices in Stage 2 of the CIRCLE lifecycle. Methods range from pre‑scripted, in‑silico tests to context‑rich, real‑world studies, shaping what forms of evidence and downstream effects can be observed. Regions illustrate classes of methods that may be selected and combined during this stage.
At the primary end of the continuum is AI benchmarking, which focuses on whether models produce “correct” or expected outputs–yielding metrics such as accuracy, efficiency, and energy use. By examining whether people can successfully navigate AI interfaces to complete tasks, methods such as usability studies begin to surface AI’s secondary effects–revealing user friction, confusion, or drop‑off points that simple benchmarks cannot detect.
Context‑sensitive methods such as red‑teaming, bias bounties, and field tests evaluate system behavior during live interactions with people under realistic or adversarial conditions [37, 30]. At the tertiary end of the continuum, longitudinal studies can be combined with other real-world signals like incident tracking to monitor cumulative downstream effects that no single snapshot can capture, such as how innovation benefits for organizations, sectors, and the broader economy evolve over time.
The “Contextual Measurement Concepts” box below lists definitions for terms related to construct development.